Question 7 1 pts The equilibrium level of real GDP is $1,000, the target level of real GDP is $1,250, and the marginal propensity to consume (MPC) is 0.80. The target can be reached if government spending is: increased by $100 billion increased by $50 billion increased by $250 billion. O held constant. Question 8 1 pts The tax multiplier equals 1 spending multiplier True False
Answer #1

Q7.Change in real GDP required = spending multiplier x change in government spending(DG)

(DG)= 1/(1 - 0.80) x ($1,250 - $1,000). DG*5=$250.

DG= $50 billion.

Hence, The target can be reached if given spending is INCREASED BY $50 billion.

Q8. False.

As, spending multiplier = 1/(1-MPC)

Whereas, tax multiplier = -MPC/(1-MPC)